# A comprehensive review of assistive technologies for children with dyslexia

### v1

@@
Abstract.
Dyslexia is a learning disability that primarily disrupts one’s ability to read, write and spell. About 15-20% of the world population is affected by dyslexia which signals it is a concerning subject. A systematic literature review was conducted on this research to evaluate trending interventions, common accessibility features, and potential barriers to existing assistive devices and to identify any research gaps. Papers published over period (2015-2023) of 8 years were reviewed and the data was meticulously analyzed. This research shows that technological-based interventions are leading interventions, especially with the use of mobile apps and augmentative realities. More innovative technologies like virtual reality, NLP, haptic, and tangible user interfaces are emerging to provide unique solutions addressing the user’s needs. Non-computing devices are less effective compared to modern digital solutions however, they provide a promising alternative.
Dyslexia, Children, Assistive Technology, Learning Disability
copyright: noneconference: ; ; copyright: none
1. Introduction
@@
2. Background
Dyslexia is a lifelong learning disorder that primarily affects reading and spelling. Since, it affects reading, some people assumes dyslexia a problem with vision. However it is a problem with language processing. Reading starts with being able to recognize and work with the sounds of spoken language. This skill is called phonological awareness. Some of the common challenges faced by children with dyslexia include:
@@ 2.4. Tactile
3. Methodology
A systematic Literature was conducted by clearly defining inclusion and exclusion criteria, specifying database sources and developing a detailed search strategy with predefined keywords. To identify the recent technological advancement publications from 2015-2023 were selected for study.
3.1. Database source
The literature review was conducted through academic search engine which includes ACM, IEE and Google Scholar. The collection of journal articles from these engines allowed for a robust search strategy, capturing the most relevant papers from multiple reputable sources within the field of study.
3.2. Keyword search
@@
3.3. Inclusion criteria
A set of criteria were defined to include research article for comprehensive review which are explained in table below.
Table 1. Inclusion Criteria
@@ 3.3. Inclusion criteria
1 Articles must be published between 2015-2024.
@@
3.4. Exclusion criteria
Exclusion criteria excludes the irrelevant journals that were out of scope to my study which were filtered out on the basis of following set of rules.
Table 2. Exclusion Criteria
@@ 3.4. Exclusion criteria
3.5. Study selection
In order to ensure the findings, directly address the research questions, inclusion and exclusion criteria were introduced. This helps me select relevant publications within the field, and it acts as the initial step in my literature review. The search string was used as a query into popular libraries including IEEE, Google Scholar and ACM. Google Scholar returned 2430 results, ACM returned 515 and lastly, IEE resulted in 181 relevant results. After filtering the articles by year of publication between 1025 to 2024, Google Scholar returned 831, ACM library returned 312 and IEEE returned 96 results which were further processed for evaluation. The final 29 papers were selected after applying all sets of defined rules from inclusion and exclusion criteria. The detailed flow diagram is presented below.
Figure 1. Selection flow diagram
4. Emerging Assistive Technologies
@@
5. Introduction to Technological Tools for Dyslexia
Technological tools are electronic, digital, or physical tools that can help people perform tasks [7]. There are wide range of technological practices designed for enhancing learning abilities for children with dyslexia. Many researchers conducted their research to resolve the learning disruption that children faced during pandemic due to COVID-19. A variety of technological devices are presented in the image below.
Figure 2. A:VR,B:Mobile app, C:Tangible user interface, D:AR
@@ 5. Introduction to Technological Tools for Dyslexia
5.1. Mobile apps
A mobile application is a software application developed specifically for use with a mobile device such as a smartphone or tablet [31].In this literature review mobile application includes all the software applications particularly tailored for mobile and tablets including augmented reality.
5.1.1. Augmented reality
Augmented reality (AR) refers to the real-time integration of digital information into a user’s environment. AR technology overlays content onto the real world, enriching a user’s perception of reality rather than replacing it [29]. The researchers Hussain et al. on designing AR games for children utilized virtual objects to enhance the reading and writing skills of children with dyslexia. The application allows children to interact with 3D alphabets and solve word puzzles, and presents increasing level of difficulty to align with their cognitive development. This innovative approach can enhance active learning among children with dyslexia.
In this study Bhatti et al. leverages augmented reality to create multimedia learning application which uses AR markers or booklets to overlay 3D objects, to aid children in comprehending different educational elements such as shapes, letters, and words—tailored for children with dyslexia. Different educational practice exercises were implemented such as recognizing sounds, understanding different animals, and their educational course work from the specially designed educational booklet that supports AR for object identification and visualizations.
Another group of researchers Fei et al. from China had integrated augmented reality with computer vision technology to facilitate learning vocabulary and provided support for writing among children with dyslexia. This application provides children to engage in identifying real-world objects through augmented reality, where children scans real world object in their environment and the application produces audio cues for targeted object.CollectiAR infact, emphasizes the role of real-world engagement, encouraging children to interact with their environment to build vocabulary skills.
In this study Lazo-Amado and Andrade-Arenas created an augmented reality (AR) application to support children with dyslexia, focusing usability and accessibility. Researchers adopted user-centric design methodology to ensure the application meets the needs of its target audience. The application incorporates tasks such as word formation, letter identifications, word identification for augmented graphics, providing engaging and interactive learning experience. While the study reported positive usability feedback, it has limitation where the user study was conducted with parents of dyslexic children rather than children.
Table 3. Descriptive categorization of Augmented reality applications for children with dyslexia.
@@ 5.1.1. Augmented reality
Research Gap Lack of engagement and interactivity in existing tools. Limited booklet design for AR. Existing word recognition and spelling training games were not able to teach students their nearby environments in real-time. Limited AR-based solutions for children with dyslexia in Latin America.
Technological Features Multiple Gaming Levels, word scramble game with timer, 3D alphabets and avatars, Rewards, Real-time alphabet identification for targeted physical object, graphics and animations, feedback support. AR-based multimedia learning using booklets as markers to overlay 3D objects, Audio cues with 3D graphics, Multiple exercises to identify letters, shapes, Numbers. Word hunt gameplay to find new words in physical environment, Audio cues to navigate target object, Provides animated feedback, Stores the captured images with words for future learning, Gameplay with spelling exercise,Tracks object and words collection for long term engagement support, Skip control Gaming Levels with 4 major activities, Completing Vowels, Selecting correct words based on image, Completing words based on image, Completing complex words without images
@@
Targeted Device smartphone and tablets. smartphones and tablets. AR-enabled smartphones. Android-based smartphones and tablets.
@@
Total Participants 21 children with age group of (7–14) years. Not specified but targeted for primary-level children. One child and two teachers in formative study. 80 (Parents)
@@
Validation Test Paired sample t-test and Wilcoxon signed-rank test To Be Evaluated Formative study with qualitative feedback from teachers and children. Expert validation on Usability, Consistency, Functionality, Integration, resulting 98% acceptance rate
@@
Findings Gamified activities such as word scramble effectively enhanced children engagement. This technology can be great help for enhanced cognitive learning via AR-based exercises. It has the potential to be a new type of supplementary vocabulary learning game for training children with dyslexia. Cognitive game improves the learning performance of children with dyslexia
@@
Gaming Activities or Tasks Word games using 3D alphabets, puzzle games with increasing difficulty levels. Letter exercises with interactive visual and audio cues using AR booklets Vocabulary scavenger hunt using computer vision to identify and spell objects in surroundings. letter recognition, and comprehension tasks using AR overlays.
@@
Software Used Unity 3D, Vuforia library for AR. Vuforia and Unity Used the computer vision model, SSD MobileNet V1. Figma, Tinkercad, MetaClassStudio.
@@
Application tailored to augmented reality can further be categorized and evaluated based on multi-sensory modality, which implements a technique, (Visual, Auditory, Kinesthetic, and tactile). Since, multi sensory model is considered as one of the optimal model enhancing reading abilities. This high level categorization provides a quick reflection and implementation details to researchers on modalities used and targeted skills.The descriptive categorization of the application is presented on table-3.
Figure 3. Multi sensory modalities used on Augmented Reality.
@@
The above table indicates that augmented reality application implements three different modalities which includes visual, auditory and kinesthetic (users movement) and represented to address mixed targeted skill sets which are crucial for enhancing the learning experience of children with dyslexia.
5.1.2. Mobile Applications
@@
The researchers Rello et al. developed a mobile application designed to address most common spelling error and word formation for children with dyslexia. The natural language processing technique was used to process the collection of handwriting from children with dyslexia to identify the most relevant words on creating the game. Most common activities include adding, removing, and rearranging letters to create correct words. The application dynamically adjusts task difficulty based on individual performance, ensuring a tailored learning experience.
@@
A group of researchers Vialatte et al. designed an application to enhance reading abilities through visual attentional training. Researchers developed symbol based visual search game where most alike symbols were presented with specific letters making identification challenging. This exercise was proven to be efficient in enhancing visual attention and reading accuracy among children with dyslexia.
@@
The researchers Jaramillo-Alcázar et al. introduced a serious game tailored to support pattern recognition and enhancing spelling abilities.On this game design children interacts with the game play and their interaction are logged into database which are sent to server for further evaluation that provides therapist holistic view on analyzing the case.
@@
The researchers Jun et al. designed game based learning system, focusing on enhancing spelling and letter identification skills. This game was designed distinguish between most alike letters (e.g., b/d and u/n) and improving overall reading abilities. Apart from this the application integrates dyslexia-friendly font and design principles to ensure accessibility and ease of use.
@@
The researchers Bigueras et al. introduces game based learning with primary focus on enhancing letter and syllable recognition skills. This application provides story based learning which make it unique among all other applications.
@@
The researchers Brennan et al. introduced mobile game with major focus on improving phonological awareness in children with dyslexia. This interactive game provides wide range of activities including exercises on developing phonological skills, such as blending of sound, identifying number of syllables, which is the foundational skills to excel reading abilities. It’s co-designed application hence, can provide more versatile learning experience.
@@
The team of researchers Gupta et al. introduced a reading assistant designed to support children with dyslexia through real-time text interaction. Utilizing optical character recognition (OCR) and text-to-speech (TTS) technologies, the system provides customizable dyslexia-friendly features such as adjustable fonts, colors, and word spacing.
@@
The following figure represents the multi-sensory modalities used on mobile applications, which highlights that mobile application heavily relies on visual and auditory cues. Furthermore mobile application addresses the demanding need on phonological awareness and reading practice among children with dyslexia.
Figure 4. Multi sensory modalities used on mobile apps.
@@
Table 4. Descriptive categorization of mobile application for children with dyslexia
Categories Dyseggxia VisioCogLetters Puzzle Piece DysPReX LaroLexia Cosmic Sounds Augmenta11y
Research Gap Paper based exercises, which were limited in enhancing user experience. Limited visual-attention training for children with dyslexia. Limited cognitive games for children with dyslexia. Phonic-based intervention methods were ineffective, DysPReX provides solution with little emphasis on phonics. Children with dyslexia have different learning capability, existing tools are less efficient. Need for phonological awareness tools in dyslexia interventions. Limited real-time, personalized reading tools for dyslexic children.
Technological Feature Six types of word exercises for enhancing learning experience. Symbol-based visual search training game, Engaged children by challenging their symbol recognition and attentional skills. Multiple games to designed to practice word games., The game sends user’s game play interaction to server to recreate more robust system . Four variety game play with interactive interface, audio cues with major focuses on visual and auditory perception. Game based learning targeted to enhance reading performance in Filipino, specifically letter recognition and syllable reading using familiar Filipino words and sight-based learning techniques. Space games for phonological awareness. OCR-based reading assistant with customizable dyslexia-friendly features.
Personalization Adaptive exercise based on user’s performance. Not specified. Can select different game levels. customizable user settings. Tailored game elements and levels based on needs. Not Specified. Adjustable fonts, colors, word spacing, and TTS integration.
Targeted Device iOS and Android devices. Not specified. PC and mobile platforms. Mobile devices and tablets. Android devices. Not Specified Not Specified.
Cross-Platform Support Yes. Not specified. No, Android only. No, Android only. No, Android only. Not specified Yes, supports iOS and Android.
@@
Total Participants 12 children with dyslexia. 21 dyslexic children. Targeted age group (5-10)years. Participant number not specified. 30 participants. 12 children, 5 teachers, and parents. 20 children with age group (9-12) years. Not Specified.
@@
Validation Methods Not Specified ANOVA, Kendall Coeff. of Concordance, Wilcoxon tests. Not Specified. Usability and satisfaction scores from participants. Pre/post-test with paired t-test for reading performance. Wilcoxon signed-rank test for pre- and post-test results. Usability feedback and performance analysis.
Findings 5000 downloads on play-store,fun, engaging, and adaptive compared to traditional exercises. Improved visual search skills. Preliminary results suggests that accessibility features and customizable difficulty levels improved user experience, making the game suitable for children with dyslexia. Positive usability feedback on user interface, functionality, learnability, and satisfaction with scores above 4.0 out of 5. Significant improvement in reading performance. Improved phonological awareness skills and engagement. Increased engagement and improved reading experience.
Gaming Activities Word games including add a letter, remove a letter,change a letter, choose the correct word ending, cut into words. Identifying alike symbols with increasing difficulties. Identifying patterns on cards, Spell the visualized object. Spelling and letter differentiation games (e.g., b/d, u/n). Story based learning, points and rewards, multi level game play, feedback, and achievement system. Interactive Storyline Theme, Open-world exploration, Mission and Progressive Level, Rewards and acheivement, Feedback and Supportive Characters, Child-Centric Design. Scan and interact with textual materials in real time.
Software Used Not specified. Not specified. Unity game engine for mobile platforms, Firebase for real-time data storage and synchronization. Unity 3D. Not specified Not specified. OCR and TTS integrated into a custom app.
5.2. Tangible User Interface

### v2

@@
Abstract
Dyslexia is a neurological learning disability that primarily disrupts one’s ability to read, write and spell, affecting an estimated 15-20% of the global population. This high prevalence underscore the importance of developing effective interventions. This study presents a systematic literature review conducted between 2015 and 2024 to evaluate current trends in assistive technologies for children with dyslexia. This research shows that technological-based interventions are leading interventions, especially with the use of mobile apps and augmentative realities. More innovative technologies like virtual reality, NLP, haptic, and tangible user interfaces are emerging to provide unique solutions addressing the user’s needs. While non-computing devices are generally less effective in comparison to modern digital solutions, they provide a promising alternative in settings with limited access to technology.
keywords:
Dyslexia, Children Assistive Technology, Learning Disability, Technology Tools
1 Introduction
@@
2 Background
Dyslexia is a lifelong learning disorder that primarily affects reading and spelling [7]. Since, it affects reading, some people assumes dyslexia a problem with vision. However it is a problem with language processing. Reading starts with being able to recognize and work with the sounds of spoken language. This skill is called phonological awareness. Some of the common challenges faced by children with dyslexia include:
@@ 2.4 Tactile
3 Methodology
A systematic Literature was conducted by clearly defining inclusion and exclusion criteria, specifying database sources and developing a detailed search strategy with predefined keywords. To identify the recent technological advancement publications from 2015-2024 were selected for study.
3.1 Database source
The literature review was conducted through academic search engine which includes ACM, IEEE and Google Scholar. The collection of journal articles from these engines allowed for a robust search strategy, capturing the most relevant papers from multiple reputable sources within the field of study.
3.2 Keyword search
@@
3.3 Inclusion criteria
A set of criteria were defined to include research article for comprehensive review which are explained in Table 1
Table 1: Inclusion Criteria
@@ 3.3 Inclusion criteria
1 Articles must be published between 2015–2024.
@@
[…]
3.4 Exclusion criteria
Exclusion criteria excludes the irrelevant journals that were out of scope to my study which were filtered out on the basis of set of rules presented in Table 2.
Table 2: Exclusion Criteria
@@ 3.4 Exclusion criteria
[…]
3.5 Study selection
In order to ensure the findings, directly address the research questions, inclusion and exclusion criteria were introduced. These criteria helped select relevant publications within the field and served as the initial step in the literature review. The search string was applied across major academic databases, including IEEE, Google Scholar, and ACM. Google Scholar initially returned 24200 results. From these, we selected top 1500 most recent publications. After screening for relevence, we excluded non-pertinent papers and finalized set of 1421 papers from this directory. In addition, ACM digital library returned 498 and IEEE Xplore resulted 151 relevent papers. Durign screening process 156 duplicates were identified across the databases, 32 papers were excluded for being non-English and 391 were excluded as they were not peer-reviewed. After filtering and applying the all above exclusion criteria the final 32 papers were selected after applying all sets of defined rules from inclusion and exclusion criteria. The detailed flow diagram is presented in Figure 1.
Figure 1: Article selection process based on inclusion and exclusion criteria. The flow diagram illustrates the screening of articles from three digital libraries (ACM, Google Scholar, and IEEE)
4 Emerging Assistive Technologies
@@
5 Introduction to Technological Tools for Dyslexia
Technological tools are electronic, digital, or physical tools that can help people perform tasks [15]. There are wide range of technological practices designed for enhancing learning abilities for children with dyslexia. Many researchers conducted their research to resolve the learning disruption that children faced during pandemic due to COVID-19. A variety of technological devices are presented in Figure 2.
Figure 2: Examples of assistive technologies used in interventions for dyslexia: (A) Virtual Reality environment, (B) Mobile learning applications, (C) Tangible user interface with illuminated letter blocks, and (D) Augmented Reality learning using 3D visualizations.
@@ 5 Introduction to Technological Tools for Dyslexia
5.1 Mobile-Assisted Learning Applications
Mobile-assisted learning applications are software applications designed to be used with mobile devices like smartphones or tablets [16]. In this literature review mobile-assisted learning application includes all the software applications particularly tailored for mobile and tablets including augmented reality.
5.1.1 Augmented reality (AR)
AR refers to the real-time integration of digital information into a user’s environment. AR technology overlays content onto the real world, enriching a user’s perception of reality rather than replacing it [17]. The researchers Hussain et al. [18] on designing AR games for children utilized virtual objects to enhance the reading and writing skills of children with dyslexia. The application allows children to interact with 3D alphabets and solve word puzzles, and presents increasing level of difficulty to align with their cognitive development. This innovative approach can enhance active learning among children with dyslexia.
In this study Bhatti et al. [19] leverages augmented reality to create multimedia learning application which uses AR markers or booklets to overlay 3D objects, to aid children in comprehending different educational elements such as shapes, letters, and words—tailored for children with dyslexia. Different educational practice exercises were implemented such as recognizing sounds, understanding different animals, and their educational course work from the specially designed educational booklet that supports AR for object identification and visualizations.
Another group of researchers Fei et al. [20] from China had integrated augmented reality with computer vision technology to facilitate learning vocabulary and provided support for writing among children with dyslexia. This application provides children to engage in identifying real-world objects through augmented reality, where children scans real world object in their environment and the application produces audio cues for targeted object. CollectiAR infact, emphasizes the role of real-world engagement, encouraging children to interact with their environment to build vocabulary skills.
In this study Amado and Arenas et al. [21] created an augmented reality (AR) application to support children with dyslexia, focusing usability and accessibility. Researchers adopted user-centric design methodology to ensure the application meets the needs of its target audience. The application incorporates tasks such as word formation, letter identifications, word identification for augmented graphics, providing engaging and interactive learning experience. While the study reported positive usability feedback, it has limitation where the user study was conducted with parents of dyslexic children rather than children.
Table 3: Descriptive categorization of Augmented Reality applications for children with dyslexia.
@@ 5.1.1 Augmented reality (AR)
Research Gap Lack of engagement and interactivity in existing tools. Limited booklet design for AR. Existing word recognition and spelling training games were not able to teach students their nearby environments in real time. Limited AR-based solutions for children with dyslexia in Latin America.
Technological Features Multiple gaming levels; word-scramble with timer; 3D alphabets and avatars; rewards; real-time alphabet identification for targeted physical objects; graphics and animations; feedback support. AR-based multimedia learning using booklets as markers to overlay 3D objects; audio cues with 3D graphics; multiple exercises to identify letters, shapes, and numbers. Word-hunt gameplay to find new words in the physical environment; audio cues to navigate to target objects; animated feedback; stores captured images with words for future learning; gameplay with spelling exercise; tracks objects/words collection for long-term engagement; skip control. Gaming levels with four activities: completing vowels; selecting correct words based on image; completing words based on image; completing complex words without images.
@@
Targeted Device Smartphones and tablets. Smartphones and tablets. AR-enabled smartphones. Android-based smartphones and tablets.
@@
Total Participants 21 children (ages 7–14). Not specified; primary-level children targeted. One child and two teachers (formative study). 80 (parents).
@@
Validation Test Paired-sample t-test and Wilcoxon signed-rank test. To be evaluated. Formative study with qualitative feedback from teachers and children. Expert validation on usability, consistency, functionality, and integration; 98% acceptance rate.
@@
Findings Gamified activities such as word scramble enhanced engagement. AR-based exercises can enhance cognitive learning. Potential supplementary vocabulary-learning game for training children with dyslexia. Cognitive game improves learning performance of children with dyslexia.
@@
Gaming Activities or Tasks Word games using 3D alphabets; puzzle games with increasing difficulty. Letter exercises with interactive visual and audio cues using AR booklets. Vocabulary scavenger hunt using computer vision to identify and spell objects in surroundings. Letter recognition and comprehension tasks using AR overlays.
@@
Software Used Unity 3D; Vuforia library for AR. Vuforia; Unity. Computer-vision model: SSD MobileNet V1. Figma; Tinkercad; MetaClassStudio.
[…]
@@
Application tailored to augmented reality can further be categorized and evaluated based on multi-sensory modality, which implements a technique (visual, auditory, kinesthetic, and tactile). Since, multi sensory model is considered as one of the optimal model enhancing reading abilities. This high level categorization provides a quick reflection and implementation details to researchers on modalities used and targeted skills. The descriptive categorization of the application is presented on Table LABEL:tab:AR_comparison_dyslexia.
Figure 3: Multi sensory modalities used on Augmented Reality.
@@
The Figure 3 indicates that augmented reality application implements three different modalities which includes visual, auditory and kinesthetic (users movement) and represented to address mixed targeted skill sets which are crucial for enhancing the learning experience of children with dyslexia.
5.1.2 Mobile Applications
@@
The researchers Rello et al. [22] developed a mobile application designed to address most common spelling error and word formation for children with dyslexia. The natural language processing technique was used to process the collection of handwriting from children with dyslexia to identify the most relevant words on creating the game. Most common activities include adding, removing, and rearranging letters to create correct words. The application dynamically adjusts task difficulty based on individual performance, ensuring a tailored learning experience.
@@
A group of researchers Vialatte et al. [23] designed an application to enhance reading abilities through visual attentional training. Researchers developed symbol based visual search game where most alike symbols were presented with specific letters making identification challenging. This exercise was proven to be efficient in enhancing visual attention and reading accuracy among children with dyslexia.
@@
The researchers Jaramillo-Alc´azar et al. [24] introduced a serious game tailored to support pattern recognition and enhancing spelling abilities.On this game design children interacts with the game play and their interaction are logged into database which are sent to server for further evaluation that provides therapist holistic view on analyzing the case.
@@
The researchers Jun et al. [25] designed game based learning system, focusing on enhancing spelling and letter identification skills. This game was designed distinguish between most alike letters (e.g., b/d and u/n) and improving overall reading abilities. Apart from this the application integrates dyslexia-friendly font and design principles to ensure accessibility and ease of use.
@@
The researchers Bigueras et al. [26] introduces game based learning with primary focus on enhancing letter and syllable recognition skills. This application provides story based learning which make it unique among all other applications.
@@
The researchers Brennan et al. [27] introduced mobile game with major focus on improving phonological awareness in children with dyslexia. This interactive game provides wide range of activities including exercises on developing phonological skills, such as blending of sound, identifying number of syllables, which is the foundational skills to excel reading abilities. It’s co-designed application hence, can provide more versatile learning experience.
@@
The team of researchers Gupta et al. [28] introduced a reading assistant designed to support children with dyslexia through real-time text interaction. Utilizing optical character recognition (OCR) and text-to-speech (TTS) technologies, the system provides customizable dyslexia-friendly features such as adjustable fonts, colors, and word spacing.
@@
The Figure 4 represents the multi-sensory modalities used on mobile applications, which highlights that mobile application heavily relies on visual and auditory cues. Furthermore mobile application addresses the demanding need on phonological awareness and reading practice among children with dyslexia. The descriptive categorization is presented in Table 4.
Figure 4: Multi sensory modalities used on mobile apps.
@@
Table 4: Descriptive categorization of mobile applications for children with dyslexia.
Categories Dyseggxia VisioCogLetters Puzzle Piece DysPReX
Research Gap Paper-based exercises limited engagement. Limited visual-attention training for dyslexic children. Few cognitive games aimed at dyslexia. Phonics-based methods ineffective; DysPReX reduces phonics focus.
Technological Features Six types of word exercises. Symbol-based visual search training; challenges symbol recognition and attention. Multiple word games; gameplay interactions logged to server. Four gameplay types; interactive UI; audio cues; focus on visual and auditory perception.
Personalization Adaptive exercises based on performance. Not specified. Selectable levels. Customizable user settings.
Targeted Device iOS and Android. Not specified. PC and mobile. Mobile and tablets.
Cross-Platform Support Yes. Not specified. No (Android only). No (Android only).
@@
Total Participants 12 children with dyslexia. 21 dyslexic children. Ages 5–10 (count not specified). 30 participants.
Validation Methods Not specified. ANOVA; Kendall’s W; Wilcoxon. Not specified. Usability and satisfaction scores.
Findings 5000+ downloads; engaging and adaptive vs. paper exercises. Improved visual search skills. Accessibility features + adjustable difficulty improved UX. Positive usability (UI, functionality, learnability, satisfaction >4/5>4/5).
Gaming Activities Add/remove/change letter; choose endings; cut into words. Identify alike symbols with increasing difficulty. Pattern recognition; spell visualized object. Spelling and letter-differentiation (e.g., b/d, u/n).
Software Used Not specified. Not specified. Unity + Firebase. Unity 3D.
[…]
Table 5: Descriptive categorization of mobile applications for children with dyslexia (continued).
Categories LaroLexia Cosmic Sounds Augmenta11y Cahyani et al. [29]
Research Gap Different learning capabilities; existing tools less efficient. Need for phonological-awareness tools in interventions. Limited real-time, personalized reading tools. Limited integration of user-centred UI with gamification and multi-sensory learning.
Technological Features Game-based reading (letter recognition, syllable reading, sight words) using familiar Filipino words. Space-themed games for phonological awareness. OCR-based reading assistant with customizable dyslexia-friendly settings. Mobile UI via design thinking; guided menus, learning paths, visual/audio/kinesthetic integration, gamified tasks, adaptive progress flow.
Personalization Tailored elements and levels. Not specified. Adjustable fonts, colours, spacing; TTS integration. Gradual content unlocking by progress.
Targeted Device Android. Not specified. Not specified. Android mobile/tablet (Figma prototype).
Cross-Platform Support No (Android only). Not specified. Yes (iOS and Android). Not specified.
Total Participants 12 children, 5 teachers, and parents. 20 children (ages 9–12). Not specified. 10 dyslexic students.
@@
Validation Methods Pre/post paired t-test for reading performance. Wilcoxon signed-rank (pre/post). Usability feedback and performance analysis. System Usability Scale: 82.
Findings Significant reading improvement. Improved phonological awareness and engagement. Increased engagement and improved reading experience. Usable, inclusive; supports structured reading/counting practice.
Gaming Activities Story-based, multilevel; feedback; rewards; achievements. Interactive storyline; missions; progressive levels; rewards; supportive characters. Scan and interact with printed text in real time. Guessing alphabets/words; ranking modes; stepwise access to activities.
Software Used Not specified. Not specified. OCR + TTS custom app. Figma.
[…]
5.2 Tangible User Interface